Find the net rate of heat transfer by radiation in watts from a skier standing in the shade, given the following. She is completely clothed in white (head to foot, including a ski mask), the clothes have an emissivity of 0.180 and a surface temperature of 12.0°C, the surroundings are at −11.5°C, and her surface area is 1.40 m2. Answer in Watts
Answer #1

Temperatures,

T1 = 12°C = 12 + 273 = 285 K

T2 = -11.5°C = -11.5 + 273 = 261.5 K

Q/t = eA σ (T14 - T24)

Q/t = 0.18 x 1.4 x 5.67 x 10^-8 (285^4 - 261.5^4)

Q/t = 27.45 W

Comment in case any doubt please rate my answer...